# Search for Long-lived Particles at Future Lepton Colliders Using Deep Learning Techniques

### v2

@@
Search for Long-lived Particles at Future Lepton Colliders Using Deep Learning Techniques
Yulei Zhang    Cen Mo    Xiang Chen School of Physics and Astronomy, Institute of Nuclear and Particle Physics, Shanghai Jiao Tong University, Shanghai 200240, China Shanghai Key Laboratory for Particle Physics and Cosmology, Key Lab for Particle Physics, Astrophysics and Cosmology (MOE), Shanghai 200240, China    Bingzhi Li    Hongyang Chen Research Center for Graph Computing, Zhejiang Lab, Hangzhou 311121, China    Jifeng Hu
hujf@m.scnu.edu.cn.
Key Laboratory of Atomic and Subatomic Structure and Quantum Control (MOE), Guangdong Basic Research Center of Excellence for Structure and Fundamental Interactions of Matter, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China Guangdong-Hong Kong Joint Laboratory of Quantum Matter, Guangdong Provincial Key Laboratory of Nuclear Science, Southern Nuclear Science Computing Center, South China Normal University, Guangzhou 510006, China
   Liang Li
liangliphy@sjtu.edu.cn.
School of Physics and Astronomy, Institute of Nuclear and Particle Physics, Shanghai Jiao Tong University, Shanghai 200240, China Shanghai Key Laboratory for Particle Physics and Cosmology, Key Lab for Particle Physics, Astrophysics and Cosmology (MOE), Shanghai 200240, China
preprint: APS/123-QED
I Introduction
In 2012, the discovery of the Higgs boson completed the final piece of the Standard Model (SM). On one hand, SM predictions match almost all experimental observations remarkably well. On the other hand, SM fails to address some key questions in the evolution of the universe, such as the existence of dark matter, the matter-antimatter imbalance puzzle and the origin of the neutrino mass. Therefore, the experimental search for new physics beyond the Standard Model (BSM) is both intriguing and necessary. So far, BSM signals remain elusive despite of numerous efforts. This typically means that either the new physics requires a higher energy threshold than the currently available experimental energy or that the coupling strength between SM and BSM particles is too small to produce a statistically significant number of observable signals. However, an alternative explanation is that BSM particles have long lifetimes and therefore can escape the detector without being detected. These BSM particles are often referred to as long-lived particles (LLPs). Recently, experimental searches for LLPs have become increasingly active. It should be noted that there are no widely-used and common definitions for LLPs and their lifetimes can vary in a wide range. In this paper we focus on LLPs produced via the rare decay of the Higgs boson, which is defined via the process $e^{+}e^{-}\to ZH(Z\to\text{inclusive},H\to X_{1}+X_{2}$). A Lepton collider can be an ideal factory for the Higgs boson, for example the proposed Circular Electron and Positron Collider (CEPC) [1, 2] aims to produce a vast amount of Higgs bosons by electron-positron annihilation at the center-of-mass energy $\sqrt{s}=240$ GeV. Fig. 1 illustrates Feynman diagrams for the LLP production process from the Higgs decay. The Higgs particle can decay into LLPs via two decay modes with two jets (type I signal) or four jets (type II signal) in the final state [42], where the neutrally charged LLPs ($X_{1}$, $X_{2}$) have identical masses and lifetimes ranging from $10^{-3}$ to $10^{2}$ nano seconds approximately. After production, LLPs can largely escape detection by the inner detector and decay into SM particles when they reach the outer detector, which can be an electromagnetic or hadronic calorimeter, or an external detector positioned at a significant distance from the interaction point.
@@ I Introduction
Figure 1: Feynman diagrams of LLP production and decay. Two Feynman diagrams are presented illustrating the generation of long-lived particles (LLPs), denoted as $X_{1}$ and $X_{2}$, through the Higgsstrahlung mechanism. On the left, the diagram shows the production of $X_{1}$ and $X_{2}$ followed by their subsequent decay into a $\nu\bar{\nu}$ pair and a $q\bar{q}$ pair, respectively, resulting in two jets. On the right, both $X_{1}$ and $X_{2}$ decay into $q\bar{q}$ pairs, leading to the four jets final state.
LLPs are sensitive probes to BSM physics, such as supersymmetry (SUSY) theory [3, 4] or dark matter models [5, 6, 7]. They can also help explain the matter-antimatter asymmetry [8, 9, 10] and the neutrino mass hierarchy [11, 12]. In experiment, LLPs often have long decay lengths and leave distinct detector signatures such as displaced vertices or jets inside detectors. Extensive experimental searches for LLPs have been made at various experiments, such as ATLAS [13], CMS [14], and LHCb [15], as well as Belle II [16], BESIII [17], Babar [18]. However, no positive discoveries have been reported so far. There could be a number of reasons for the absence of LLPs signals, low signal-over-background ratio and complexity in event reconstruction with displaced objects are two major ones. In this paper we present a novel method utilizing deep learning techniques to overcome these hurdles. Such techniques can be applied directly to low-level detector information without the need to fully reconstruct the whole event. A deep learning based algorithm has been developed to learn comprehensive LLPs topologies and kinematic features that can distinguish LLPs from various SM backgrounds with high efficiencies and extremely low fake rates. The algorithm has been trained with full simulation Monte Carlo (MC) samples corresponding to an integrated luminosity of 20.0 $20.0\text{~{}ab}^{-1}$, including LLPs signal samples generated using the event generator MadGraph 3.0.1 [19], and background samples corresponding to $e^{+}e^{-}\to q\bar{q}$, $e^{+}e^{-}\to Z/W$ and $e^{+}e^{-}\to ZH\to\text{inclusive}$ processes generated using the event generator Whizard 1.95 [20, 21]. Other minor backgrounds such as pileup background and cosmic ray background have also been studied and found to be negligible in this analysis.
II Method
General Analysis Strategy
After taking into detector effects, we apply a clustering algorithm to the energy, position, and time of the detector hits. Subsequently, we adopt an ML-based analysis strategy by feeding the clustered hits information into advanced neural networks for multi-label classification. All hit energies have been calibrated using MC samples fully simulated under CEPC conceptual design [24, 34, 35]. The energy resolution for the electromagnetic calorimeter is at $16\%/\sqrt{E/\text{GeV}}\oplus 1\%$, and for the hadronic calorimeter and muon detector it is at $60\%/\sqrt{E/\text{GeV}}\oplus 1\%$. The hit time resolution is set to be 1 nanosecond. Since the ML-based analyses cluster hits at a coarse scale of several centimeters, the finer hit position resolution provided by the detector becomes less critical because the clustering process effectively averages out the hit positions over larger areas.
We define two classifications for the dominating SM backgrounds: the 2-fermion background accounting for $e^{+}e^{-}\rightarrow q\bar{q}$ and $e^{+}e^{-}\to ZH\to\text{inclusive}$ processes, and the 4-fermion background including $e^{+}e^{-}\rightarrow W/Z$ and $e^{+}e^{-}\to ZH\to\text{inclusive}$ processes. We define three classifications for the LLPs signal according to the number of detectable LLPs  [43]: 0, 1, 2. After training with MC simulations, we apply the trained neural networks to statistically independent MC samples to obtain five classification outputs. These outputs serve as inputs of the XGBoost algorithm [36] to further enhance the signal over background discrimination. A final selection is placed on the XGBoost discriminant output to reject all backgrounds and calculate the signal yield.
After the discriminant selection, we employ a toy model to generate 10000 pseudo-experiments based on the predicted signal yields under the zero background condition. We then build a test statistics with the likelihood ratio method [37] to determine 95% C.L. upper limits on $\mathcal{B}(H\rightarrow\text{LLPs})$. A parameter $\epsilon_{V}$ is defined as the ratio $\frac{BR(X\rightarrow\nu\bar{\nu})}{BR(X\rightarrow q\bar{q})}$. $\epsilon_{V}$ can be treated as a fixed parameter or floating parameter. The latter represents a more model-independent approach.
Convolutional Neural Network
@@ II Method
For the CNN approach, each event with raw detector hits is transformed into a two-channel image with $200\times 200$ pixels in $(R,\phi)$, where $R$ represents the distance of the position of the hits to the interaction point ($0-6$ m), and $\phi$ represents the azimuthal angle of the hits ($0-2\pi$). The first channel (energy channel) represents a sum of all hits energies in a pixel. The second channel (time channel) represents the time difference ($\Delta t$) associated with the most energetic hit in a pixel as shown in Eq. 1.
$\Delta t=t_{\text{hit},\text{maxE}}-r_{\text{hit},\text{maxE}}/c$ (1)
@@
$t_{\text{hit},\text{maxE}}$ is the time of the hit with the maximum energy in a pixel, $r_{\text{hit},\text{maxE}}$ is the Euclidean distance from the interaction point to hit location, and $c$ is the speed of light in vacuum. A selection of hits energy larger than $0.1$ GeV is placed to suppress contributions from extremely low energy hits of secondary particles. Figure 2 shows images of hits for signal and background events. A larger circle indicates a higher energy deposition and a darker color represents a smaller time difference. The event images illustrate the discriminating power of CNN to separate LLPs signals with displaced energies or objects from SM backgrounds. By feeding event images to ResNet18 neural networks [22] shown in Fig. 3, a multi-label classification is performed. The CNN model is trained using 5 GPUs with a batch size of 256 per GPU. The learning rate is adaptively set, beginning at $10^{-4}$ and progressively decreasing to $10^{-6.5}$. The training employs the Adam optimizer with a weight decay of 0.01. The training duration is set at 10 epochs, and the model’s final configuration is determined by the epoch that achieves the lowest validation loss. The distributions of the training and testing losses for CNN are shown in Sec. 1.3 in ”Supplementary Information”.
@@
Figure 2: Event visualization of CNN. CNN event images for the Type-I signal (left) and the Type-II signal (right) with backgrounds. Circles represent detector hits in the calorimeter and squares represent detector hits inside the tracker. Darker pixels represent hits with smaller time differences, and bigger pixels represent hits with larger energy. The decay vertex of LLPs is marked with a star symbol.
@@
Graph Neural Network
For the GNN approach, the raw calorimeter hits and tracker hits belonging to an event are clustered and fed into a heterogeneous event graph with both the calorimeter-type and tracker-type of nodes. The nodes of the same detector type are fully connected to construct edges. To perform the hits clustering in the calorimeter, we first identify the most energetic calorimeter hit and group it together with its neighboring hits that are within a spatial distance of $50~{}mm$. We require the number of neighboring hits larger than 3 and the total number of hits in a cluster larger than 10. We then iterate this procedure for all hits in an event to construct the calorimeter-type nodes. All hits failing the requirements are not clustered and excluded from the GNN. The momentum $p$ of each calorimeter hit is defined as parallel to its position: $p_{i}=\frac{i}{r}E$, where $i$ represents the position along the $x$, $y$, or $z$ axis, with $z$ as the beam line direction, $x$ and $y$ in the transverse plane, and $y$ pointing towards the ground. $r$ is the radial distance $r=\sqrt{x^{2}+y^{2}+z^{2}}$, and $E$ is the hit’s energy. The tracker hits are binned into $5~{}\times~{}6$ blocks based on their R-$\phi$ positions. In each block, the tracker hits are clustered into a tracker-type node taking an arithmetic average of all hits positions. The definitions of nodes and edges for the GNN event graph are summarized in Table 1.

### v5

@@
Abstract
Long-lived particles (LLPs) provide an unambiguous signal for physics beyond the Standard Model (BSM). They have a distinct detector signature, with decay lengths corresponding to lifetimes of around nanoseconds or longer. Lepton colliders allow LLP searches to be conducted in a clean environment, and such searches can reach their full physics potential when combined with machine learning (ML) techniques. This experimental study, utilizing comprehensive full simulation data samples, focuses on LLP searches resulting from Higgs decay in $e^{+}e^{-}\to ZH$. We demonstrate that, by employing deep neural network approaches the LLP signal efficiency can be improved up to 95% for an LLP mass around 50 GeV and a lifetime of approximately 1 nanosecond, while rejecting all SM backgrounds. Furthermore, the signal sensitivity for the branching ratio of Higgs decaying into LLPs reaches a state-of-art limit of $9.7\times 10^{-7}$ with a statistics of $4\times 10^{6}$ Higgs.
footnotetext: $\diamond$ These authors contributed equally to this work.footnotetext: * Authors to whom any correspondence should be addressed.
1 Introduction
In 2012, the discovery of the Higgs boson completed the final piece of the Standard Model (SM) [1, 2]. Remarkably, SM predictions align with almost all experimental observations. However, the SM does not address several critical questions in the universe’s evolution, such as the existence of dark matter [3, 4, 5], the matter-antimatter imbalance puzzle [6, 7] and the origin of the neutrino mass [8, 9]. Consequently, the search for new physics beyond the Standard Model (BSM) is both intriguing and necessary. Despite numerous efforts, BSM signals have remained elusive, suggesting that new physics may either require a higher energy threshold than currently achievable or that the coupling strength between SM and BSM particles is too weak to produce a statistically significant number of observable signals. Amidst this scenario, the hypothesis that BSM particles could possess long lifetimes, evading detection, has transitioned from a nascent idea to a widely accepted and vigorously pursued avenue within the physics community. These particles, often called long-lived particles (LLPs), serve as sensitive probes into BSM physics, such as supersymmetry (SUSY) theory [10, 11] or dark matter models [12]. LLPs could also provide insights for the matter-antimatter asymmetry [13] and the neutrino mass hierarchy, with their lifetimes varying widely. Recently, the search for LLPs has intensified across several experiments, such as ATLAS [14, 15], CMS [16, 17], and LHCb [18], as well as Belle II [19], BESIII [20], Babar [21], marking a keen interest in their potential to reveal new aspects of physics, even though definitive discoveries are yet to be made.
LLPs are characterized by their extended decay lengths and distinctive detector signatures, such as displaced vertices or internal jets, allowing them to evade detection by the inner detectors after being produced. Instead, they decay into SM particles upon reaching the outer detectors, which could be electromagnetic or hadronic calorimeters, or even external detectors positioned significantly away from the interaction point. The elusive nature of LLP signals can be attributed to factors such as a low signal-over-background ratio and the complexity of event reconstruction involving displaced objects. Consequently, specialized reconstruction techniques are required to effectively detect and measure LLPs.
Searching for LLPs using machine-learning (ML) assisted approaches at hadron colliders [22, 23, 24] has emerged as a promising strategy to surmount these experimental challenges. Such research employs ML techniques for object-level tagging to identify LLP signatures following comprehensive event reconstruction. The efficacy of this tagging is dependent on a constrained set of high-level input variables. The demanding environment at hadron colliders complicates background control, possibly necessitating neural networks with higher granularity and more extensive input features. Furthermore, the deployment of advanced triggers powered by rapid neural networks is crucial for the efficient detection of LLP-like events at hadron colliders [25, 26].
@@ 1 Introduction
Figure 1: Feynman diagrams of LLP production and decay. Two Feynman diagrams are presented illustrating the generation of long-lived particles (LLPs), denoted as $X$, through the Higgsstrahlung mechanism. On the left, the diagram shows the production of $XX$ followed by their subsequent decay into a $\nu\bar{\nu}$ pair and a $q\bar{q}$ pair, respectively, resulting in two jets. On the right, both $X$ decay into $q\bar{q}$ pairs, leading to the four jets final state.
Electron-positron colliders, such as the Circular Electron Positron Collider (CEPC) [27] and the International Linear Collider (ILC) [28], provide clean collision environments with precise initial-state conditions. The Higgs boson production via $e^{+}e^{-}\rightarrow ZH$ at these facilities offers a clean channel to explore rare LLP decays, benefiting from well-defined initial kinematics and significantly reduced backgrounds compared to hadron colliders like the LHC [29, 30]. The primary goal of this analysis is to measure the cross section of LLP production through Higgs boson decays at future lepton colliders. Given that the SM Higgs production cross section at electron-positron colliders is accurately calculable and well-known, the LLP production cross section can be directly derived by multiplying the predicted SM $ZH$ cross section by the assumed branching ratios (BRs) of Higgs into LLPs. Therefore, these branching ratios are the fundamental parameters that we aim to constrain in this analysis.
The Higgs boson can decay into a pair of long-lived particles (LLP, denoted as $X$), each of which subsequently decays into final state objects such as jets, charged leptons, or neutrinos. In the scope of our study, the search strategy is inspired by the glueball and Gauge Mediated Supersymmetry Breaking (GMSB) models [31, 32], where such LLPs naturally emerge and have accessible decay modes as defined by the process of $H\rightarrow XX\rightarrow\textrm{jets}$. This focus stems from the distinguishable detector signatures that the jet decay products can offer.
Figure 1 shows Feynman diagrams for the LLP production process from the Higgs decay in lepton colliders with two jets (type I signal) or four jets (type II signal) in the final state. We focus on scenarios where the mixing between the new scalar particle $X$ and the Higgs boson is negligible. This assumption allows us to isolate and analyze specific decay channels of $X$ without the complexities introduced by universal couplings to all SM particles that would result from significant $X$-Higgs mixing. While $X$ could, in principle, decay into gauge bosons, our study focuses on the hadronic and invisible decay channels. This focus is motivated by scenarios such as effective $Z^{\prime}$ models, where the new scalar $X$ couples preferentially to quarks and neutrinos over charged leptons and gauge bosons [33, 34, 35]. The invisible decay channel of $X$ is defined to include any final states that result in missing energy signatures, such as neutrinos or potential dark matter candidates. This approach aligns with various BSM scenarios where $X$ serves as a mediator between the SM and the dark sector, leading to missing energy signatures in the detector. Additionally, hadronic and invisible final states pose significant experimental challenges due to large QCD backgrounds and missing energy signatures, respectively, making them compelling channels for detailed study.
Other potential decay channels such as $H\rightarrow XX\rightarrow\textrm{invisible}$ and $H\rightarrow XX\rightarrow\textrm{leptons}$ are not considered in this study. Particularly, the $H\rightarrow XX\rightarrow\textrm{invisible}$ channel, classified as a type III signal, results in decay products that leave no detectable signature within the detector’s sensitivity range. Such invisibly decaying LLPs do not contribute to observable events that can be analyzed within our current experimental framework. Consequently, these decays are more appropriately addressed in analyses specifically dedicated to $H\rightarrow\textrm{invisible}$ decays [36]. Similarly, while $H\rightarrow XX\rightarrow\textrm{leptons}$ presents an interesting decay channel, the current study prioritizes jet signatures due to their robust detection capabilities and the analytical focus on hadronic final states. Future research may expand to encompass these additional decay modes, taking into account different event topologies and analysis methodologies.
In this study, we investigate the feasibility and sensitivity of employing advanced machine learning methods, particularly deep learning techniques such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), to learn the intricate topologies and kinematic features of LLPs. We conduct a comparative analysis of the two distinct machine learning architectures to evaluate their respective performances in reconstructing LLP signals. Each model is independently trained and assessed, providing insights into their individual capabilities and limitations within the context of LLP detection. By applying these methods directly to low-level detector data, we achieve remarkable improvements in LLPs signal reconstruction efficiency and background suppression. The search sensitivity for LLPs at future $e^{+}e^{-}$ colliders achieved using deep learning techniques significantly surpasses that of traditional methods, establishing state-of-the-art limits on the Higgs branching ratio into LLPs. This approach enables us to derive stringent constraints on the branching ratios, thus enhancing our ability to identify potential LLP signals and probe BSM physics effectively. Importantly, the ML-based methodology we present offers broad applicability, promising an easy adaptation for future lepton collider experiments, including the CEPC [27], ILC [28], FCC [37], and CLIC [38].
The remainder of this paper is structured as follows: details on the samples used are provided in Section “Signal and Background Samples”. In Section “Method”, we describe the methodology of our analysis, including analysis strategy, reconstruction algorithms, and neural network architectures. Sections “Results” and “Discussion” present detailed numerical results and performance evaluations. Section “Conclusion” summarizes our findings and discusses implications for future LLP searches.
2 Signal and Background Samples
In this study, we utilize fully simulated Monte Carlo (MC) samples for both the signal and background analyses. To generate the long-lived particle (LLP) signal samples, MadGraph 3.0.1 [39] is employed to cover both type I and type II signals across various mass points (50 GeV, 10 GeV, and 1 GeV) and lifetimes (0.001 ns, 0.1 ns, 1 ns, 10 ns, 100 ns). A total of $1.5\times 10^{7}$ events are produced, distributing $10^{6}$ events for each combination of mass point and lifetime.
The backgrounds predominantly consist of SM processes resulting in jets in the final state, notably $e^{+}e^{-}\rightarrow\bar{q}q$ and $e^{+}e^{-}\rightarrow VV$, where $VV$ includes both $WW$ and $ZZ$. Additionally, the SM process $e^{+}e^{-}\rightarrow ZH\rightarrow\textrm{inclusive}$ is also considered, given its significance as the main production mode for SM Higgs. These SM processes are simulated using Whizard 1.95 [40, 41]. A total of $10^{7}$ events is simulated for both the $e^{+}e^{-}\rightarrow\bar{q}q$ and $e^{+}e^{-}\rightarrow VV$ processes, while the $e^{+}e^{-}\rightarrow ZH\rightarrow\textrm{inclusive}$ process has $10^{6}$ events simulated. The top quark and tri-boson related backgrounds have negligible cross sections at a center-of-mass energy of 240 GeV and therefore are not considered. Minor backgrounds such as pileup background and cosmic ray background are also neglected in this analysis (Appendix B). Details on the cross-sections and simulated statistics for the LLP signal and SM background are provided in Table 1.
For each Monte Carlo (MC) sample, a full detector simulation is performed following the CEPC conceptual design [42]. The CEPC detector comprises a silicon-based vertexing and tracking system, a Time Projection Chamber (TPC) tracker, a high-granularity calorimetry system, a 3 Tesla superconducting solenoid magnet, and a muon detection system integrated within the magnet’s iron return yoke. The calorimetry system consists of an electromagnetic calorimeter (ECAL) and an iron-scintillator hadronic calorimeter (HCAL). Simulations are carried out using MOKKAC software [43], a dedicated Geant4-based tool customized for the CEPC framework, which accounts for comprehensive particle showering effects and calibrated detector responses. The calorimeter resolutions are modeled to achieve an energy resolution of $16\%/\sqrt{E/\textrm{GeV}}\oplus 1\%$ for the ECAL and $60\%/\sqrt{E/\textrm{GeV}}\oplus 1\%$ for the HCAL and muon detector. In addition, detector hits are assigned a time resolution of 1 nanosecond. Following full detector simulation, no further event or object reconstruction is applied; rather, the digitized detector hits are directly utilized as inputs for the subsequent ML-based analyses.
Table 1: Cross sections and simulated statistics of the signal and background samples.
Process LLP signal SM ZH $q\bar{q}$ $ZZ$ $WW$
$\sigma[\mathrm{fb}]$ 203.66 54106.86 1110.37 16721.77
Statistics $1.5\times 10^{7}$ $1.0\times 10^{6}$ $1.0\times 10^{7}$ $1.1\times 10^{6}$ $8.9\times 10^{6}$
3 Method
3.1 General Analysis Strategy
We develop a novel analysis strategy leveraging deep learning techniques based on low-level detector information, circumventing the need for event or object reconstruction. This approach marks a significant departure from traditional selection-based methods.
Initially, a signal acceptance selection is applied to ensure that at least one LLP is within the geometric boundaries of the CEPC detector [42], with acceptance factors detailed in Table 2. Subsequently, signal events are classified based on the number of LLP decays detectable within the detector and their consequent final states. This classification yields three categories: $\mathcal{D}_{0}$, $\mathcal{D}_{1}$, and $\mathcal{D}_{2}$, denoting events with 0, 1, and 2 detectable LLPs leading to jet final states, respectively.
Acceptance (%) Lifetime [ns]
Mass [GeV] 0.001 0.1 1 10 100
For the background analysis, events are categorized based on their final-state jet multiplicities into two categories: 2-jet and 4-jet. The dominant background for the 2-jet category arises from the process $e^{+}e^{-}\rightarrow q\bar{q}$ , while the 4-jet background primarily originates from the diboson production process $e^{+}e^{-}\rightarrow VV$. Additionally, the inclusive SM Higgs production process ($e^{+}e^{-}\rightarrow ZH\rightarrow$ inclusive) contributes to both the 2-jet and 4-jet categories.
Figure 2: Workflow of event classification for LLP signals. This chart outlines the event classification process utilizing CNN and GNN models, followed by XGBoost analysis. The process starts with formatting detector hits for NN input, progresses through 5-class classification, and concludes with an XGBoost-enhanced selection to differentiate the signal from the SM background, finally determining signal efficiency.
Unlike the conventional method of event and object reconstruction followed by complex kinematic and object selection, our study adopts a direct application of two neural network (NN) models: a Convolutional Neural Network (CNN) and a Graph Neural Network (GNN), for the purpose of event classification and performance evaluation. Our method involves transforming low-level detector information on the energy, position, and time of detector hits, into formats suitable for these neural models, leading to a comprehensive classification of events into five distinct categories, as detailed in Figure 2:
  • LLPs signals: $\mathcal{D}_{0}$, $\mathcal{D}_{1}$, $\mathcal{D}_{2}$
  • SM backgrounds: 2-jet, 4-jet
The classification output is a set of prediction scores, indicating the probability of an event belonging to each of the five categories. Following this NN-based classification, an XGBoost algorithm [44] is implemented to use these prediction scores as inputs to generate a single output score. This score discriminates between signal events and SM background.
To achieve a background-free analysis, we apply stringent event selection criteria based on XGBoost output scores to reject all SM background events. Events resulting in 0 detectable LLPs ($\mathcal{D}_{0}$), despite being classified as signal events during the neural network training stage to capture the full range of LLP behaviors, are subsequently excluded from the final signal extraction and limit calculation, since such events produce no experimentally distinguishable signatures. The fractions of events with exactly one ($\mathcal{D}_{1}$) or two ($\mathcal{D}_{2}$) detectable LLPs are derived for both type I and type II signals using Monte Carlo samples. The overall signal efficiency reported in this analysis thus reflects the combined selection efficiency of events containing at least one detectable LLP. Finally, based on the achieved signal efficiencies, we estimate the signal sensitivity and derive exclusion limits on the Higgs branching ratio into LLPs. We employ two independent neural networks, CNN and GNN, to conduct a comparative study, evaluating their respective strengths in LLP reconstruction and signal identification.
3.2 Convolutional Neural Network
@@
For the CNN approach, each event with low-level detector information is transformed into a two-channel image with $200\times 200$ pixels in $(R,\phi)$, where $R$ represents the distance of the position of the hits to the interaction point ($0-6$ m) and $\phi$ represents the azimuthal angle of the hits ($0-2\pi$). The first channel (energy channel) represents the sum of all hits energies in a pixel. The second channel (time channel) represents the time difference ($\Delta t$) associated with the most energetic hit in a pixel as shown in Eq. 1.
$\Delta t=t_{hit,maxE}-{r_{hit,maxE}}/c$ (1)
@@
$t_{hit,maxE}$ is the time of the hit with the maximum energy in a pixel, $r_{hit,maxE}$ is the Euclidean distance from the interaction point to the hit location, and $c$ is the speed of light in vacuum. A selection of hits energy larger than $0.1$ GeV is placed to suppress contributions from extremely low energy hits of secondary particles. Figure 3 shows images of hits for signal and background events. The event images illustrate the discriminating power of CNN to separate LLP signals with displaced energies or objects from SM backgrounds. By feeding event images to ResNet18 neural networks [45] shown in Figure 4, a multi-label classification is performed. The CNN model is trained using 5 GPUs with a batch size of 256 per GPU. The learning rate is adaptively set, beginning at $10^{-4}$ and progressively decreasing to $10^{-6.5}$. The training employs the Adam optimizer with a weight decay of 0.01. The training duration is set at 10 epochs, and the model’s final configuration is determined by the epoch that achieves the lowest validation loss. The distributions of the training and testing losses for CNN are shown in Appendix C.
@@
Figure 3: Event visualization of CNN. Left: an event with one detectable LLP ($\mathcal{D}_{1}$, Type-I signal) featuring a decay vertex at [4.5 m, 2.3 rad]; Right: an event with two detectable LLPs ($\mathcal{D}_{2}$, Type-II signal) with decay vertices at [3.69 m, 2.16 rad] and [0.99 m, 4.37 rad]. Circles represent detector hits in the calorimeter and squares represent detector hits inside the tracker. Darker pixels represent hits with smaller time differences, and bigger pixels represent hits with larger energy. Note that the varying circle sizes are solely for visualization purposes; in the analysis, the pixel size remains fixed. The decay vertex of LLPs is marked with a star symbol.
@@
3.3 Graph Neural Network
In the GNN approach [46], low-level calorimeter and tracker hits from an event are clustered and then integrated into a heterogeneous event graph, encompassing both calorimeter-type and tracker-type nodes. We implement a simple clustering method based on spatial proximity, grouping hits around the most energetic ones within a predefined region. This method is intentionally chosen for its computational simplicity, aiming to quickly and effectively reduce data dimensionality without requiring extensive event reconstruction. Given that lepton collider environments are relatively clean compared to hadron colliders, more complex clustering algorithms offer limited improvements in performance but at increased computational cost. Sensitivity studies confirmed that varying clustering parameters, such as minimum-hit thresholds and cluster radii, had negligible impact on signal efficiency and background rejection efficiency, thereby validating the robustness and suitability of our approach. The clustering process helps to minimize the complexity of the graph and significantly reduces memory usage during training. Furthermore, nodes of the same detector type are comprehensively interconnected, facilitating the formation of edges within the graph. This structure optimizes the analysis by ensuring efficient data processing and interaction modeling between different components of the detector hits. For the clustering of hits in the calorimeter, a point cloud composition method based on energy screening and distance thresholding is introduced. The most energetic calorimeter hit is first identified as the core hit. Subsequently, hits within a predefined distance threshold of $50~{}mm$ from the core hit are merged, constructing a graphical node structure. It is required that the number of merged hits exceeds 3 and the total number of hits in a cluster exceeds 10. This procedure is then iterated for all hits in an event to construct the calorimeter-type nodes. Hits that fail to meet these requirements are not clustered and are excluded from the GNN. The momentum $p$ of each calorimeter hit is defined as parallel to its position: $p_{i}=\frac{i}{r}E$, where $i$ represents the position along the $x$, $y$, or $z$ axis, with $z$ as the beam line direction, $x$ and $y$ in the transverse plane, and $y$ pointing towards the ground. $r$ is the radial distance $r=\sqrt{x^{2}+y^{2}+z^{2}}$, and $E$ is the hit’s energy. The tracker hits are binned into $5~{}\times~{}6$ blocks based on their R-$\phi$ positions. In each block, the tracker hits are clustered into a tracker-type node taking an arithmetic average of all hit positions. The definitions of nodes and edges for the GNN event graph are summarized in Table 3.